How many moles of CaSO4 are present in 12.0 grams of CaSO4?
1
Expert's answer
2020-10-28T08:43:06-0400

Number of moles= mass/molar mass

Molar mass of CaSO4=[40+32+(16×4)]= 40+32+64= 136g

Number of moles= 12/136= 0.088 mole